OpenStudy (anonymous):

Given a polynomial f(x), if (x + 3) is a factor, what else must be true? f(3) = 0 f(-3) = 0 f(0) = 3 f(0) = -3

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

if \(x+3)\) is a factor then it must look like \[f(x)=(x+3)(something )\] and so \[f(-3)=(-3+3)(something)=0\times something =0\]